OpenStudy (anonymous):

Which transformation describes the equation from its parent equation? 1/3 f(x)

OpenStudy (anonymous):

y - 4 = 2(x + 3) 2(y - 3) = x + 4 y + 3 = 3(x - 4) 2(y + 4) = x + 3

OpenStudy (anonymous):

@jim_thompson5910